# Dust destruction by the supernova remnant forward shock in a turbulent interstellar medium

### v1

@@ 2.2 SN explosion and SNR evolution
The simulation output of Sect. 2.1 is used as the starting point for our MHD SNR simulations. In these simulations, we set off an SN with an explosion energy of 1051 erg using the SN feedback method of Arepo that injects thermal energy into 40 central neighboring Voronoi cells, corresponding to a radius of 0.08–0.21 pc in our simulations. The SNR evolution is followed for 10 kyr. The final snapshots are shown in Fig. 3 (see also Appendix A for the full evolution of the simulations, including movies). In addition to the mass refinement of Arepo, we implemented refinement with respect to the pressure gradient to follow the SNR forward shock in greater detail, since it is the key region where dust destruction happens. Cell splitting and merging is enabled, so that the total amount of cells almost doubles throughout the simulation as the SNR grows. Thus, the average distance between neighboring cell centers at the end of the simulations is of the order of 0.030 and 0.060 pc for the 20 and 40 pc box size, respectively.
@@
In the SNR simulations, radiative cooling is considered. We assume collisional ionization equilibrium and electron-ion equipartition as we do not follow electron and ion temperatures separately, or ionization states in detail. The used cooling implementation in the code depends on whether the gas has been shocked, which is determined by its temperature. Unshocked ISM gas has an initial temperature of 104 K and cools according to detailed molecular and atomic cooling functions, assuming solar metallicity and a simplified treatment of gas-phase molecular chemistry (Glover2012; 2016Bertram). High density filaments cool down efficiently before they are being shocked whereas voids remain at higher temperatures due to their low density. Shocked gas is identified in cells where the temperature exceeds 104 K. We assume that this gas mixes with SN ejecta, with the latter dominating the cooling process. Since our simulations do not include explicitly injected ejecta, the shocked ISM effectively takes on the role of the ejecta. To emulate the conditions in the ejecta of Cassiopeia A (Cas A), which we use as a reference, we adopt oxygen-dominated cooling based on the CHIANTI cooling function, following the approach of Kirchschlager2019. This method was chosen because applying the ISM cooling prescription to the ejecta resulted in a faster expansion of the SNR, due to insufficient cooling, leading to a divergence from the observed evolution of SNRs similar to Cas A. This approximation is considered reliable only during the early stages of the SN, when the shocked mass remains comparable to the expected ejecta mass. Because the swept-up ISM material is heated beyond the ISM cooling threshold, it is cooled according to the ejecta prescription, resulting in a slight overestimation of cooling at the shock front, particularly at late times in the high density simulations.
Thermal conduction is not considered. Its timescale can be calculated by (e.g. Orlando2005)
@@
where $l$ is the length scale of interest and $\kappa(T)=5.6\times 10^{-7}T^{-\frac{5}{2}}\,\mathrm{erg\,s^{-1}\,K^{-1}\,cm^{-1}}$ is the thermal conductivity. The shortest timescale of $\sim$13 kyr is reached at the contact of the thin shell and the SNR interior of the high ISM density simulations. However, the forward shock is expanding on shorter timescales in our simulations so that thermal conduction can be neglected. Although at very early times of a few hundreds of years, the thermal conduction timescale between the forward shock and the ISM is on the order of only a few thousand years, the ISM is heated more rapidly by the expansion of the forward shock than by conduction at this early stage.
Having set all initial parameters, the SN is set off in the driven turbulent ISM box. Because stars usually excavate their surroundings with stellar winds and radiative feedback, most of them explode in a low density cavity (e.g. Lamers1999; Rogers2013; Fichtner2024). Modeling the stellar feedback mechanisms in detail is out of the scope for this paper, however, we mimic the impact of pre-SN feedback by always placing the SN in a void of the turbulence, also ensuring that all simulations are easily comparable. The void is chosen manually by shifting the 3D turbulence box to a position with a large void in the center. Therefore, the forward shock encounters filaments only after several hundred to a few thousand years of evolution for the higher Mach number simulations as is also the case for SNe that explode in a wind-driven bubble. The voids in our strong turbulence simulations are roughly 10–20 pc large, which is comparable to the observed and predicted wind-driven cavity sizes of 10–100 pc of massive stars, strongly dependent on the stellar mass and ISM density (McCray1983; Georgy2013). If the SNe were to explode closer to dense filamentary structures instead of voids, the filaments would be encountered at higher velocity, resulting in more efficient dust destruction in the early SNR stages. This is discussed in Appendix B.
After evolving the SNR for 10 kyr in MHD simulations, we study the dust dynamics and destruction in post-processing simulations using Paperboats since Arepo does not include detailed dust processing. These dust simulations are run from snapshots of the SNR simulation, meaning that their time steps are larger than the hydrodynamic time steps. Since the forward shock is faster in the beginning and slows down significantly (from $v>10^{4}\,\mathrm{km\,s^{-1}}$ to $v<400\,\mathrm{km\,s^{-1}}$), the first few hundred years of evolution must have a higher temporal resolution than the last few hundred years. Thus, we took 512 snapshots of the MHD simulations at logarithmically spaced time steps to feed the post-processing simulations, resulting in time steps of 0.15–120 yr for the dust simulations.
@@ 2.3 Dust dynamics
We post-process the 512 snapshots from each MHD simulation with the Paperboats code (Kirchschlager2019) in order to determine the dust dynamics. This code was developed to study the dust destruction of the reverse shock in dense ejecta clumps (Kirchschlager2019; Kirchschlager2020; Kirchschlager2023; Kirchschlager2024b) and later also used for forward shock dust destruction (Kirchschlager2022; Kirchschlager2024a). It simulates the dust by introducing a certain dust mass according to an initially fixed gas-to-dust mass ratio parameter that we here adopt to be 100, which is close to the value assumed in dust models characterizing the diffuse ISM in the Milky Way (e.g. Draine2007; Compiegne2011; Jones2013). Based on the amount of dust, its properties, and the fluid conditions from the MHD simulations, Paperboats can determine how the dust will be affected by the gas and other dust grains via gas and plasma drag, Lorentz acceleration, thermal and kinetic sputtering, effects of grain-grain collisions (fragmentation and vaporization), and accretion in the following MHD snapshots. The rotational disruption of dust grains by mechanical torques (METD) or radiation (RATD, Hoang2020) is not included in Paperboats. These effects were recently evaluated in the context of SNR dust destruction (Martinez-Gonzales2025), and were found to only have a minor impact on the overall dust destruction. However, this could change for our model setup and will be part of analysis in the near future.
@@
Initially, the dust is given the same velocity as the gas. As such, dust in the central cells would not be given an initial velocity so that it would not move out of the center. Therefore, we do not start with the first snapshot of the MHD simulation but from snapshot 10, which relates to only a few years after the explosion. The dust is then added with the dust grain-size distribution of Mathis, Rumpl and Nordsieck (MRN; Mathis1977) from 5 to 250 nm, thus a power law distribution $a^{-\gamma}$ with $\gamma=3.5$. To account for grain destruction and accretion, the grain sizes in our simulations range from 0.6 to 350 nm, which are separated into 20 logarithmically spaced bins. In total, we perform 4 dust destruction simulations for each SNR simulation, considering only sputtering, or sputtering and grain-grain collisions to assess the impact of grain-grain collisions, and either silicate or carbonaceous dust to study how the dust destruction efficiency varies for different grain species. More detailed information about the code implementation of Paperboats can be found in Kirchschlager2019; Kirchschlager2020; Kirchschlager2023.
@@
The gas-dust interactions simulated by Paperboats assume that the ISM gas consists of a single atomic species. We model the gas as being dominated by hydrogen atoms and therefore use a mean molecular weight of $\mu=1$ for our dust simulations. As a result, hydrogen atoms must also dominate in the region where ISM gas and ejecta gas mix. This approximation becomes increasingly accurate as the SNR evolves since the mass of ISM gas in the SNR interior increases over time. In our simulations, the remnant encounters at least 150 M of ISM material during 10 kyr, which is significantly more than typical SNR ejecta masses (e.g. $\sim$4 M for Cas A). Thus, hydrogen dominates the total mass in the SNR interior for most of the evolution, although the cooling is still dominated by oxygen. For consistency, we also use this simplified molecular weight for our density calculations in both the MHD and dust simulations. However, solar abundances are generally assumed for the rest of the MHD calculations.
@@
Since Paperboats does not support a moving mesh yet, we convert the Voronoi grid of Arepo to a Cartesian grid that can be used by Paperboats. The Cartesian grid has 900 cells in each coordinate direction, almost twice as high as the initial average distance of cell centers in the MHD simulations. The dust simulation resolution was found to be optimal for our setup and to not induce significant errors during the simulations, as also discussed in Appendix C. During the conversion from Voronoi to Cartesian grid, the total mass of the system changes due to the rearrangement of the cells. However, our chosen resolution was high enough to only change the mass by 0.2%, which has a minor impact on the dust destruction.
@@
Furthermore, we only study the central 2D layer because a full 3D dust calculation would have been numerically too expensive. The final amount of Cartesian grid cells is therefore $900\times 900\times 1$ cells in $x$-, $y$-, and $z$-direction. To ensure that the investigated layer is representative of the whole 3D SNR, we performed the dust simulations on three different perpendicular planes. The dust destruction of the forward shock in these planes can alter by several per cent as shown in Appendix D, depending on the different detailed density distribution in the slice. Thus, dust destruction values in this paper always refer to the average of these three planes, unless otherwise stated.
@@
To evaluate the dust destruction efficiency of only the encountered ISM dust instead of the dust in the full simulation box, we had to trace the region which is shocked by the SNR in all snapshots. Since the SNR is still hot and fast after 10 kyr, it was possible to follow it by using a temperature or velocity gradient threshold, which were implemented into Paperboats. The temperature threshold worked best for the subsonic and homogeneous cases and traces all cells which are or were above the temperature of $1.1\times 10^{4}\,\mathrm{K}$ in any of the MHD snapshots. This threshold was chosen to correctly follow the SNR in all time steps, even in the first few where the ISM is at $10^{4}\,\mathrm{K}$ but cools down rapidly. For the trans- and supersonic simulations, voids do not cool quickly but can be compressed and thus heated by the slight diffusion of the filaments. Therefore, the temperature threshold was increased to $3\times 10^{4}\,\mathrm{K}$ for the trans- and supersonic simulations with a low ISM density. For the intermediate and high density cases, however, it is possible that, at a given snapshot, shocked gas cells cool down to temperatures lower than the ones retained in the unshocked gas in voids. Hence, we used a $v\Delta v_{i}$ threshold that traces all cells that are or were above it at any given time step for the trans- and supersonic simulations of these cases. $\Delta v_{i}$ is here the velocity difference of neighboring cells in the coordinate direction $i\in\{x,y,z\}$. Because this threshold is resolution dependent and the forward shock evolves differently depending on the ISM density and Mach number, we adjusted the $v\Delta v_{i}$ threshold for each of the four simulations for which it was used. This was done by carefully fine-tuning the threshold until no ISM gas cells but all SNR forward shock cells are traced. Since we track all cells that are or once were above this threshold, we finally follow the whole SNR interior over time. We manually checked that the SNR evolution is followed correctly for all of our snapshots. Thereby, we could follow the dust destruction and grain-size distributions of the SNR shocked region more easily, noting that mixed cells at the shock front may be traced too early or too late. However, cells at the boundary are only a small fraction of all traced SNR cells so that the impact on the total dust destruction rate is negligible.
@@ 3 Results
In this section, we will discuss the impact of the different initial ISM densities and turbulent Mach numbers on the SNR evolution, and analyze how they influence the dust destruction efficiency, considering different dust species and destruction mechanisms. The gas density distribution of all SNR simulations at 10 kyr is shown in Fig. 3 (see Appendix A for the full evolution, including movies). Depending on the considered parameter set, the total dust destruction lies between 27–92%, corresponding to 0.85–11.0 M of destroyed dust in the first 10 kyr of SNR evolution. The given destroyed dust masses are not the dust masses in a 2D plane but an extrapolation of the 2D destruction rates multiplied with the 3D encountered dust mass. The dust destruction results, averaged over the three investigated orthogonal planes, are compared in Table 1 and the results of the $z=0$ plane dust simulations are visualized in Fig. 4 where the rows show the different Mach numbers ($\mathcal{M}=0,0.3,1,3$) and the columns show the average ISM densities ($n=1,10,100\,\mathrm{cm^{-3}}$).
Table 1: Overview of the results after 10 kyr for all dust simulations carried out in this paper. [^2] The simulations vary in their considered average ISM density $n$, turbulent Mach number $\mathcal{M}$, and dust species. The values in the top 12 lines belong to the simulations that consider grain-grain collisions and sputtering while the bottom 12 lines belong to the sputtering-only simulations. The values are averaged over three orthogonal 2D planes. Their standard deviation from the mean was used as the error of the mean. The average end shock velocity $v_{\text{shock}}^{\text{end}}$ was estimated by locating the SNR forward shock from a momentum threshold. The uncertainty in this value is due to the use of different thresholds, all of which trace the shock reasonably well.
@@
Figure 4: Dust destruction efficiencies (left $y$-axis) and masses (right $y$-axis) in the $z=0$ slice. This destruction efficiency refers to the ratio between the destroyed dust mass at each time step to the total encountered dust mass at the last considered snapshot. Different Mach numbers of $\mathcal{M}=0,0.3,1,\text{ and }3$, and average ISM densities of $n=1,10,\text{ and }100\,\mathrm{cm^{-3}}$ are shown. For each Mach number and ISM density parameter set, 4 different dust simulations were carried out, including either silicate (Sil) or carbonaceous (Car) dust, and either grain-grain collisions and sputtering (gg+sp) or only sputtering (sp).
3.1 Low density ISM
In a low density environment ($n=1\,\mathrm{cm^{-3}}$, left column of Fig. 3), the forward shock is not efficiently slowed down by ISM gas (247–388 km s-1 at 10 kyr, Table 1) and can thus expand to large sizes, reaching up to 24–40 pc, depending on the studied turbulent Mach number. Due to the low amount of shocked gas in this ISM and the fixed initial gas-to-dust mass ratio, small amounts of dust are encountered, resulting in low destroyed dust masses of 1.4-1.5 M for silicate dust, and 0.85–0.97 M for carbonaceous dust after 10 kyr in our simulations. The destruction efficiencies are also rather low with 43–57% for silicate and 27–36% for carbonaceous dust (also shown in Table 1 and Fig. 4), which may be unexpected given that the shock front is fast and hot throughout the whole simulation, implying efficient thermal and kinetic sputtering. However, sputtering of dust grains can only occur when they collide with the hot and fast ions of the shock front. Since these ions have a rather low number density and move through a low density ISM, which has an even lower dust grain number density of 0.01 cm-3 on average, the energetic ions of the forward shock do not collide often with dust grains. Thus, they cannot transfer their energy to the ISM grains (i.e. sputter them) efficiently. Due to the low grain number density, grain-grain collisions rarely occur, leading to differences of maximum 2.6% compared to the sputtering-only simulations. Therefore, the processed grain-size distributions with and without grain-grain collisions are similar for grains larger than 5 nm, showing a survival of most grains above $\sim$40 nm and significant destruction fractions for smaller grains. Many grains smaller than 5 nm are generated, increasing even more with included grain-grain collisions that produce many small fragments (see Appendix E).

### v2

@@ 2.2 SN explosion and SNR evolution
The simulation output of Sect. 2.1 was used as the starting point for our MHD SNR simulations. In these simulations, we set off an SN with an explosion energy of 1051 erg using the SN feedback method of Arepo that injects thermal energy into 40 central neighboring Voronoi cells, corresponding to a radius of 0.08–0.21 pc in our simulations. The SNR evolution was followed for 10 kyr. We show the final snapshots in Fig. 3 (see also Appendix A for the full evolution of the simulations, including movies). In addition to the mass refinement of Arepo, we implemented refinement with respect to the pressure gradient to follow the SNR forward shock in greater detail, since it is the key region where dust destruction happens. Cell splitting and merging were enabled, so that the total number of cells almost doubled throughout the simulation because the SNR grew. Thus, the average distance between neighboring cell centers at the end of the simulations was of the order of 0.030 and 0.060 pc for the 20 and 40 pc box size, respectively.
@@
In the SNR simulations, radiative cooling is considered. We assumed collisional ionization equilibrium and electron-ion equipartition as we do not follow electron and ion temperatures separately, or ionization states in detail. The used cooling implementation in the code depends on whether the gas has been shocked, which was determined by its temperature. Unshocked ISM gas had an initial temperature of 104 K and cooled according to detailed molecular and atomic cooling functions, assuming solar metallicity and a simplified treatment of gas-phase molecular chemistry (Glover and Clark 2012; Bertram et al. 2016). High-density filaments cool down efficiently before they are shocked, whereas voids remain at higher temperatures, due to their low density. We identified shocked gas in cells where the temperature exceeded 104 K. We assumed that this gas mixed with SN ejecta, with the latter dominating the cooling process. Since our simulations did not include explicitly injected ejecta, the shocked ISM effectively took on the role of the ejecta. To emulate the conditions in the ejecta of Cassiopeia A (Cas A), which we used as a reference, we adopted oxygen-dominated cooling based on the CHIANTI cooling function (Dere et al. 1997; Del Zanna et al. 2015), following the approach of Kirchschlager et al. (2019). This method was chosen because applying the ISM cooling prescription to the ejecta resulted in a faster expansion of the SNR, due to insufficient cooling, leading to a divergence from the observed evolution of SNRs similar to Cas A. This approximation is considered reliable only during the early stages of the SN, when the shocked mass remains comparable to the expected ejecta mass. Because the swept-up ISM material is heated beyond the ISM cooling threshold, it was cooled according to the ejecta prescription, resulting in a slight overestimation of cooling at the shock front, particularly at late times in the high-density simulations.
We did not consider thermal conduction. Its timescale can be calculated by (e.g., Orlando et al. 2005)
@@
where $l$ is the length scale of interest and $\kappa(T)=5.6\times 10^{-7}T^{-\frac{5}{2}}\,\mathrm{erg\,s^{-1}\,K^{-1}\,cm^{-1}}$ is the thermal conductivity. The shortest timescale of $\sim$13 kyr is reached at the contact of the thin shell and the SNR interior of the high ISM density simulations. However, the forward shock is expanding on shorter timescales in our simulations so that thermal conduction can be neglected. Although at very early times of a few hundred years, the thermal conduction timescale between the forward shock and the ISM is on the order of only a few thousand years, the ISM is heated more rapidly by the expansion of the forward shock than by conduction at this early stage.
Having set all initial parameters, the SN was set off in the driven turbulent ISM box. Because stars usually excavate their surroundings with stellar winds and radiative feedback, most of them explode in a low-density cavity (e.g., Lamers and Cassinelli 1999; Rogers and Pittard 2013; Fichtner et al. 2024). Modeling the stellar feedback mechanisms in detail is beyond the scope of this paper; however, we mimicked the impact of pre-SN feedback by always placing the SN in a void of the turbulence, also ensuring that all simulations are easily comparable. The void was manually chosen by shifting the 3D turbulence box to a position with a large void in the center. Therefore, the forward shock encountered filaments only after several hundred to a few thousand years of evolution for the higher Mach number simulations as is also the case for SNe that explode in a wind-driven bubble. The voids in our strong turbulence simulations were roughly 10–20 pc large, which is comparable to the observed and predicted wind-driven cavity sizes of 10–100 pc of massive stars, strongly dependent on the stellar mass and ISM density (McCray 1983; Georgy et al. 2013). If the SNe were to explode closer to dense filamentary structures instead of voids, the filaments would be encountered at higher velocity, resulting in more efficient dust destruction in the early SNR stages. This is discussed in Appendix B.
After evolving the SNR for 10 kyr in MHD simulations, we studied the dust dynamics and destruction in post-processing simulations using Paperboats since Arepo does not include detailed dust processing. These dust simulations are run from snapshots of the SNR simulation, meaning that their time steps are larger than the hydrodynamic time steps. Since the forward shock is faster in the beginning and slows down significantly (from $v>10^{4}\,\mathrm{km\,s^{-1}}$ to $v<400\,\mathrm{km\,s^{-1}}$), the first few hundred years of evolution must have a higher temporal resolution than the last few hundred years. Thus, we took 512 snapshots of the MHD simulations at logarithmically spaced time steps to feed the post-processing simulations, resulting in time steps of 0.15–120 yr for the dust simulations.
@@ 2.3 Dust dynamics
We post-processed the 512 snapshots from each MHD simulation with the Paperboats code (Kirchschlager et al. 2019) in order to determine the dust dynamics. This code was developed to study the dust destruction of the reverse shock in dense ejecta clumps (Kirchschlager et al. 2019, 2020, 2023, 2024b) and was later also used for forward shock dust destruction (Kirchschlager et al. 2022, 2024a). It simulates the dust by introducing a certain dust mass according to an initially fixed gas-to-dust mass ratio parameter that we adopted to be 100, which is close to the value assumed in dust models characterizing the diffuse ISM in the Milky Way (e.g., Draine and Li 2007; Compiègne et al. 2011; Jones et al. 2013). Based on the amount of dust, its properties, and the fluid conditions from the MHD simulations, Paperboats can determine how the dust will be affected by the gas and other dust grains via gas and plasma drag, Lorentz acceleration, thermal and kinetic sputtering, effects of grain-grain collisions (fragmentation and vaporization), and accretion in the following MHD snapshots. The rotational disruption of dust grains by mechanical torques (METD) or radiation (RATD, Hoang and Lee 2020) is not included in Paperboats. These effects were recently evaluated in the context of SNR dust destruction (Martínez-González 2025), and were found to only have a minor impact on the overall dust destruction. However, this could change for our model setup and will be part of analysis in the near future.
@@
Initially, the dust was given the same velocity as the gas. As such, dust in the central cells would not be given an initial velocity so that it would not move out of the center. Therefore, we did not start with the first snapshot of the MHD simulation but from snapshot 10, which relates to only a few years after the explosion. The dust was then added with the dust grain-size distribution of Mathis, Rumpl and Nordsieck (MRN; Mathis et al. 1977) from 5 to 250 nm, and thus a power-law distribution $a^{-\gamma}$ with $\gamma=3.5$. To account for grain destruction and accretion, the grain sizes in our simulations ranged from 0.6 to 350 nm, which were separated into 20 logarithmically spaced bins. In total, we performed 4 dust destruction simulations for each SNR simulation, considering only sputtering, or sputtering and grain-grain collisions to assess the impact of grain-grain collisions, and either silicate or carbonaceous dust to study how the dust destruction efficiency varies for different grain species. More detailed information about the code implementation of Paperboats can be found in Kirchschlager et al. (2019, 2020, 2023).
@@
The gas-dust interactions simulated by Paperboats assume that the ISM gas consists of a single atomic species. We modeled the gas as being dominated by hydrogen atoms and therefore used a mean molecular weight of $\mu=1$ for our dust simulations. As a result, hydrogen atoms must also dominate in the region where ISM gas and ejecta gas mix. This approximation becomes increasingly accurate as the SNR evolves since the mass of ISM gas in the SNR interior increases over time. In our simulations, the remnant encountered at least 150 M of ISM material during 10 kyr, which is significantly more than typical SNR ejecta masses (e.g., $\sim$4 M for Cas A). Thus, hydrogen dominated the total mass in the SNR interior for most of the evolution, although the cooling was still dominated by oxygen. For consistency, we used this simplified molecular weight for our density calculations in both the MHD and dust simulations. However, solar abundances were generally assumed for the rest of the MHD calculations.
@@
Since Paperboats does not support a moving mesh yet, we converted the Voronoi grid of Arepo to a Cartesian grid that could be used by Paperboats. The Cartesian grid had 900 cells in each coordinate direction, almost twice as high as the initial average distance of cell centers in the MHD simulations. The dust simulation resolution was found to be optimal for our setup and to not induce significant errors during the simulations, as also discussed in Appendix C. During the conversion from Voronoi to Cartesian grid, the total mass of the system changed due to the rearrangement of the cells. However, our chosen resolution was high enough to only change the mass by 0.2%, which had a minor impact on the dust destruction.
@@
Furthermore, we only studied the central 2D layer because a full 3D dust calculation would have been numerically too expensive. The final amount of Cartesian grid cells was therefore $900\times 900\times 1$ cells in $x$-, $y$-, and $z$-direction. To ensure that the investigated layer was representative of the whole 3D SNR, we performed the dust simulations on three different perpendicular planes. The dust destruction of the forward shock in these planes can differ by several per cent as shown in Appendix D, depending on the different detailed density distribution in the slice. Thus, dust destruction values in this paper always refer to the average of these three planes, unless otherwise stated.
@@
To evaluate the dust destruction efficiency of only the encountered ISM dust instead of the dust in the full simulation box, we had to trace the region which was shocked by the SNR in all snapshots. Since the SNR is still hot and fast after 10 kyr, it was possible to follow it by using a temperature or velocity gradient threshold, which were implemented into Paperboats. The temperature threshold worked best for the subsonic and homogeneous cases and traced all cells which are or were above the temperature of $1.1\times 10^{4}\,\mathrm{K}$ in any of the MHD snapshots. This threshold was chosen to correctly follow the SNR in all time steps, even in the first few where the ISM is at $10^{4}\,\mathrm{K}$ but cools down rapidly. For the trans- and supersonic simulations, voids did not cool quickly but can be compressed and thus heated by the slight diffusion of the filaments. Therefore, we increased the temperature threshold to $3\times 10^{4}\,\mathrm{K}$ for the trans- and supersonic simulations with a low ISM density. For the intermediate- and high-density cases, however, it is possible that, at a given snapshot, shocked gas cells cool down to temperatures lower than the ones retained in the unshocked gas in voids. Hence, we used a $v\Delta v_{i}$ threshold that traced all cells that are or were above it at any given time step for the trans- and supersonic simulations of these cases. This $\Delta v_{i}$ threshold is the velocity difference of neighboring cells in the coordinate direction $i\in\{x,y,z\}$. Because this threshold is resolution dependent and the forward shock evolves differently depending on the ISM density and Mach number, we adjusted the $v\Delta v_{i}$ threshold for each of the four simulations for which it was used. This was done by carefully fine-tuning the threshold until no ISM gas cells but all SNR forward shock cells were traced. Since we tracked all cells that are or once were above this threshold, we finally followed the whole SNR interior over time. We manually checked that the SNR evolution was followed correctly for all of our snapshots. This meant that we could follow the dust destruction and grain-size distributions of the SNR shocked region more easily, noting that mixed cells at the shock front may be traced too early or too late. However, cells at the boundary were only a small fraction of all traced SNR cells so that the impact on the total dust destruction rate is negligible.
@@ 3 Results
In this section we discuss the impact of the different initial ISM densities and turbulent Mach numbers on the SNR evolution, and analyze how they influence the dust destruction efficiency, considering different dust species and destruction mechanisms. The gas density distribution of all SNR simulations at 10 kyr is shown in Fig. 3 (see Appendix A for the full evolution, including movies). Depending on the considered parameter set, the total dust destruction lies between 27–92%, corresponding to 0.85–11.0 M of destroyed dust in the first 10 kyr of SNR evolution. The given destroyed dust masses are not the dust masses in a 2D plane but an extrapolation of the 2D destruction rates multiplied with the 3D encountered dust mass. The dust destruction results, averaged over the three investigated orthogonal planes, are compared in Table 1 and the results of the $z=0$ plane dust simulations are visualized in Fig. 4 where the rows show the different Mach numbers ($\mathcal{M}=0,0.3,1,3$) and the columns show the average ISM densities ($n=1,10,100\,\mathrm{cm^{-3}}$).
Table 1: Overview of the results after 10 kyr for all dust simulations carried out in this paper. [^2] The simulations vary in their considered average ISM density $n$, turbulent Mach number $\mathcal{M}$, and dust species. The values in the top 12 lines belong to the simulations that consider grain-grain collisions and sputtering, while the bottom 12 lines belong to the sputtering-only simulations. The values are averaged over three orthogonal 2D planes. Their standard deviation from the mean was used as the error of the mean. The average end shock velocity $v_{\text{shock}}^{\text{end}}$ was estimated by locating the SNR forward shock from a momentum threshold. The uncertainty in this value is due to the use of different thresholds, all of which trace the shock reasonably well.
@@
Figure 4: Dust destruction efficiencies (left $y$-axis) and masses (right $y$-axis) in the $z=0$ slice. This destruction efficiency refers to the ratio of the destroyed dust mass at each time step to the total encountered dust mass at the last considered snapshot. Different Mach numbers of $\mathcal{M}=0,0.3,1,\text{ and }3$, and average ISM densities of $n=1,10,\text{ and }100\,\mathrm{cm^{-3}}$ are shown. For each Mach number and ISM density parameter set, four different dust simulations were carried out, including either silicate (Sil) or carbonaceous (Car) dust, and either grain-grain collisions and sputtering (gg+sp) or only sputtering (sp).
3.1 Low-density ISM
In a low-density environment ($n=1\,\mathrm{cm^{-3}}$, left column of Fig. 3), the forward shock is not efficiently slowed down by ISM gas (247–388 km s-1 at 10 kyr, Table 1) and can thus expand to large sizes, reaching up to 24–40 pc, depending on the studied turbulent Mach number. Due to the low amount of shocked gas in this ISM and the fixed initial gas-to-dust mass ratio, small amounts of dust are encountered, resulting in low destroyed dust masses of 1.4-1.5 M for silicate dust, and 0.85–0.97 M for carbonaceous dust after 10 kyr in our simulations. The destruction efficiencies are also rather low with 43–57% for silicate and 27–36% for carbonaceous dust (also shown in Table 1 and Fig. 4), which may be unexpected given that the shock front is fast and hot throughout the whole simulation, implying efficient thermal and kinetic sputtering. However, sputtering of dust grains can only occur when they collide with the hot and fast ions of the shock front. Since these ions have a rather low number density and move through a low-density ISM, which has an even lower dust grain number density of 0.01 cm-3 on average, the energetic ions of the forward shock do not collide often with dust grains. Thus, they cannot transfer their energy to the ISM grains (i.e., sputter them) efficiently. Due to the low grain number density, grain-grain collisions rarely occur, leading to differences of maximum 2.6% compared to the sputtering-only simulations. Therefore, the processed grain-size distributions with and without grain-grain collisions are similar for grains larger than 5 nm, showing a survival of most grains above $\sim$40 nm and significant destruction fractions for smaller grains. Many grains smaller than 5 nm are generated, increasing even more with included grain-grain collisions that produce many small fragments (see Appendix E).